OpenStudy (anonymous):

Use 5 step problem solving method. Marilyn's plan for a rectangular goat pen is too small. The perimeter needs to be more than 118 feet. If the width needs to be nine feet shorter than the length, what should the length be? *Again, tips to help with these types of problems would be AWESOME!

OpenStudy (anonymous):

P = 2w + 2l P > 118 l = ? w = l - 9 2 (l - 9) + 2l > 118 2l - 18 + 2l > 118 4l - 18 > 118 4l > 100 l > 25

OpenStudy (anonymous):

Your 3rd to last line...I don't understand where the +18 went that you had to do to the 118? Wouldn't it be L=36?

OpenStudy (anonymous):

I mean, shouldn't it be: 118 + 18 would be 136 = 136/4 = 34 ?

OpenStudy (anonymous):

118 is not < than 82...? I'm CONFUSED??

OpenStudy (anonymous):

ok, I'm going with L>=34 Thank you for helping me figure out how to put this into an equation! :)

OpenStudy (anonymous):

Sorry. You're right. I should have added 18 to both sides. Good catch.